# Identification and estimation of mediational effects of longitudinal modified treatment policies

### v1

@@ 3.3 Estimation
The foregoing analysis allows us to construct estimators of the parameter $\theta$ given in Theorem 1 with desirable statistical properties such as asymptotic normality and efficiency. The theory of semiparametric inference allows for the analysis of estimators of parameters (e.g., causal effects) without restricting $\mathsf{P}$ to a parametric model. Standard references include van der Vaart (1998) and Bickel et al. (1997). In our case, we are particularly interested in nonparametric inference which places no restrictions on $\mathsf{P}$ other than the identification assumptions, particularly those related to positivity.
@@
$\mathsf{D}_{L,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{{}^{\prime}}_{t+1,s}\mathsf{H}_{t,s}\{\mathsf{Q}_{Z,s+1}-\mathsf{Q}_{L,s}\}+\sum_{s=t+1}^{\tau}\mathsf{K}^{{}^{\prime}}_{t+1,s}\mathsf{H}_{t,s-1}\{\mathsf{Q}_{L,s}-\mathsf{Q}_{Z,s}\}+\mathsf{Q}_{L,t}$ (5)
$\mathsf{D}_{Z,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{{}^{\prime}}_{t,s}\mathsf{H}_{t,s}\{\mathsf{Q}_{Z,s+1}-\mathsf{Q}_{L,s}\}+\sum_{s=t}^{\tau}\mathsf{K}^{{}^{\prime}}_{t,s}\mathsf{H}_{t,s-1}\{\mathsf{Q}_{L,s}-\mathsf{Q}_{Z,s}\}+\mathsf{Q}_{Z,t}$ (6)
$\mathsf{D}_{M,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{*}_{t,s}[\prod_{k=t}^{s-1}\mathds{1}(M_{k}=m_{k})][\mathds{1}(M_{s}=m_{s})\mathsf{Q}_{M,s+1}-Q_{M,s}]+Q_{M,t}$ (7)
@@
The EIF for $\theta$ in the nonparametric model is given by $\mathsf{S}(X,\eta)=\sum_{\bar{m}\in\bar{\mathcal{M}}}\left[\{\mathsf{D}_{Z,1}(X,\bar{m};\eta)-\varphi(\bar{m})\}\lambda(\bar{m})+\{\mathsf{D}_{M,1}(X,\bar{m};\eta)-\lambda(\bar{m})\}\varphi(\bar{m})\right]$.
@@
The algorithm used to estimate $\theta$ is given as Algorithm 1 in Díaz et al. (2023b), where in our case the IPW weights are replaced with probability ratios as described in Section 3.2. The algorithm consists of the following steps. First, we set $\hat{\mathsf{D}}_{Z,\tau+1}=Y$ and $\hat{\mathsf{D}}_{M,\tau+1}=1$. Then, beginning at $t=\tau$, we calculate $\hat{\mathsf{Q}}_{L,t}$ as the nonparametric regression prediction of $\hat{\mathsf{D}}_{Z,t+1}$, onto preceding variables. Then, we nonparametrically estimate the functions $\mathsf{g}$ and $\mathsf{G}$ (as in Section 3.2), and $\hat{\mathsf{D}}_{L,t}$ is calculated by plugging in all the nuisance functions estimated so far into its definition (in equation 5). Similarly, $\hat{\mathsf{Q}}_{Z,t}$ comes from a nonparametric regression of $\hat{\mathsf{D}}_{L,t}$, and $\hat{\mathsf{Q}}_{M,t}$ comes from a nonparametric regression of $\mathbbm{I}(M_{t}=m_{t})\hat{\mathsf{D}}_{M,t+1}$. Then $\hat{\mathsf{D}}_{Z,t}$ is estimated from equation 6 and $\hat{\mathsf{D}}_{M,t}$ is estimated from equation 7. This is repeated for $t=\tau-1,\tau-2,\dots,1$. Finally, the mean of $\hat{\mathsf{D}}_{Z,1}$ is used as an estimator for $\phi$, and the mean of $\hat{\mathsf{D}}_{M,1}$ is taken as an estimator of $\lambda$, which are plugged into the identifying formula of Theorem 1 to estimate $\theta$. (Note that the use of the $\mathsf{D}$-functions in this way is motivated by Lemma 1 of Díaz et al. (2023b).) The estimated variance of the efficient influence function (Theorem 2) is used to estimate the standard error. The procedure makes use of cross-fitting in all regression functions to ensure that the asymptotic linearity result above holds without restrictions on the complexity of the form of the machine learning estimators, thereby yielding valid statistical inference under mild conditions (Zheng and van der Laan, 2011; Chernozhukov et al., 2018).
Our estimation algorithm has been implemented in the open-source R package lcmmtp (Williams and Díaz, 2023).
4 Motivating application
4.1 Background
In the early weeks of the COVID-19 pandemic, recommendations from international health organizations advocated for early intubation of patients in an effort to safeguard healthcare workers from contracting infection and to reduce complications resulting from late intubations when the patient is unresponsive or apneic (Papoutsi et al., 2021). Over time, however, as clinicians became more familiar with the progression of the disease, guidance evolved toward postponing intubation, partially driven by reports showing high mortality rates for mechanically ventilated patients, potentially attributable to heightened risk of secondary infections, lung injury, and damage to other organs including the kidneys, due to ventilation (Bavishi et al., 2021).
[…]
This data was previously described and studied in Díaz et al. (2023a), which treated AKI and other-cause mortality as competing risks; we take a different perspective by considering AKI as a mediator in the pathway from IMV to mortality. This orientation presents a more intricate view of the causal connections among these variables. Essentially, AKI and mortality are not only competing risks; they can jointly occur, with AKI potentially influencing mortality. By examining AKI’s mediating role, we aim to unravel the indirect ways in which IMV may contribute to mortality via AKI. Should this causal pathway exist, it could provide novel opportunities for medical interventions aimed at preventing or treating AKI, subsequently reducing the mortality rate. (As mentioned in Section 1, these options include the modulation of intravascular volume status Grams et al. (2011); Glassford and Bellomo (2011) and various pharmacological interventions (Pickkers et al., 2022).) Further, we analyze how various baseline variables are associated with the estimated direct and indirect effects and discuss how such considerations might inform treatment decisions regarding IMV.
We note that this data, inasmuch as it includes treatment decisions made in the face of considerable uncertainty and unstable guidelines regarding a novel disease, is especially valuable from the point of view of causal inference for observational data. This is because causal inference relies on the assumptions of positivity (that individuals at any covariate level have positive probability of being assigned to either treatment condition) and ignorability (that the treatment-outcome relationship is not confounded by unmeasured variables). If treatment decisions are made based on a generally agreed-upon function of clinical indicators, then one or the other of these conditions is likely to be violated; if all the clinical indicators are included in the data, then positivity is doubtful; if some of the clinical indicators are omitted, then ignorability is doubtful. By contrast, in the current context, there can be enough variation in treatment decision-making to preserve positivity even with many important measured covariates. Practical violations of positivity can be diagnosed from observed data by looking at whether there is overlap in the covariate distributions corresponding to different treatment options (Zhu et al., 2021). Lack of ignorability cannot generally be diagnosed by observed data alone, but “confounding by indication,” where non-adjustment variables influence both treatment indication and outcomes, has been noted as a common issue in observational studies of medical treatments (Kyriacou and Lewis, 2016). For example, observational studies that link selective serotonin reuptake inhibitors to suicidal side effects must account for the fact that patients prescribed treatment for depression are already at a higher risk for such outcomes (Lee and Chang, 2022).
This study uses data at the daily level, and study time begins on the day of hospitalization. The treatment variable is categorized into three levels: no supplemental oxygen, supplemental oxygen not including IMV, and IMV. The main goal is to estimate the overall causal effect of IMV on mortality/survival rates and to decompose it into effects operating through AKI and effects operating independently of it. To measure the total effect of IMV on survival, we posit an intervention that would delay IMV by one day among patients who received IMV for more than one day and would prevent IMV for patients who received IMV for exactly one day. We refer to this intervention as a “one-day delay in intubation.” Formally, for $A_{t}\in\{0,1,2\}$ corresponding to the end of day $t$, 00 indicates no supplemental oxygen, $1$ indicates supplemental oxygen not including IMV, and $2$ indicates IMV. We consider the following intervention:
$\mathbbm{d}_{t}(a_{t},h_{t})=\begin{cases}1&\text{ if }a_{t}=2\text{ and }a_{s}\leq 1\text{ for all }s<t,\\ a_{t}&\text{ otherwise.}\end{cases}$ (8)
@@ 4.2 Time-varying variables, mediator, and modified treatment policy
The structure of the data is as follows. Baseline confounders include age, sex, race, ethnicity, body mass index (BMI), hospital location, home oxygen status, and comorbidities (e.g., hypertension, history of stroke, diabetes mellitus, coronary artery disease, active cancer, cirrhosis, asthma, chronic obstructive pulmonary disease, interstitial lung disease (ILD), HIV infection, and immunosuppression), and are included in $L_{1}$. Time-dependent confounders $Z_{t}$ and $L_{t}$ include vital signs (e.g., highest and lowest respiratory rate, oxygen saturation, temperature, heart rate, and blood pressure), laboratory results (e.g., C-Reactive Protein, BUN creatinine ratio, activated partial thromboplastin time, creatinine, lymphocytes, neutrophils, bilirubin, platelets, D-dimer, glucose, arterial partial pressure of oxygen, and arterial partial pressure of carbon dioxide). In cases of missing baseline confounders, mean substitution with an additional variable to indicate missingness is used. For missing data at later timepoints, the last observation is carried forward. Patients are censored at their day of hospital discharge, as AKI and vital status were unknown after this point.
In constructing the analytical dataset, it is important that the variables preserve the time-ordering of the DAG in Figure 1. Therefore, we constructed the dataset using the following procedure. First, we categorize patients into three groups: those whose first event was an intubation, those whose first event was AKI, and those who had neither event. For patients in the first group, we anchor the creation of their record on the time at which the intubation event occurred, denoted $t_{I}$. The variables $L_{t_{I-1}}$ and $Z_{t_{I-1}}$ are then recorded as those observed in approximately 24-hour windows before $t_{I}$, where $L_{t_{I-1}}$ contains variables measured in the first half of this approximate 24-hour interval, and $Z_{t_{I-1}}$ contains variables measured in the second half of this approximate 24-hour window. The time interval is approximate because the difference in hours between hospital admission and intubation is not necessarily a multiple of 24. The mediator $M_{t_{I-1}}$ is then set as 00, and the exposure $A_{t_{I-1}}$ is recorded as the respiratory support in the previous 24-hour window. This process is iterated until we have data for all the timepoints prior to $t_{I}$. For times greater than or equal to $t_{I}$, we follow a similar approach, but in addition to capturing $Z_{t}$, $L_{t}$, and $A_{t}$, we measure whether the patient was diagnosed with AKI at every time window. AKI was defined using creatinine values in accordance with the ’Kidney Disease: Improving Global Outcomes’ definition (KDIGO) definition (Khwaja, 2012). Either of the following criteria was required: (a) serum creatinine change of greater than or equal to 0.3 mg/dL within 48 h, or (b) serum creatinine greater than or equal to 1.5 times the baseline serum creatinine known or assumed to have occurred within the past 7 days.
If and when a patient is diagnosed with AKI (call this time $t_{M}$), we begin a new 24-hour interval, i.e. $L_{t_{M}}$, to ensure that the temporality of variables measured after $t_{M}$ is maintained. A similar procedure was used for patients whose first event was the diagnosis of AKI. Data for patients in the third group, who had neither event, was simply divided into approximate 24-hour time intervals using the anchor times of hospital admission and censoring or death date. During this process, we ensured that measurement of $L_{t}$ precedes $Z_{t}$ within each time interval. A GitHub repository of analysis code is available at https://github.com/CI-NYC/lcmmtp-application.
4.3 Results
We used our proposed methods to estimate the total, direct, and indirect effects on mortality of invasive mechanical ventilation through acute kidney failure. All regressions were estimated using lcmmtp (Williams and Díaz, 2023) and Super Learner (Van der Laan et al., 2007), building a convex combination of three predictors: generalized linear models with $\ell_{1}$ penalty, multivariate adaptive regression splines, and extreme gradient boosting of regression trees. To speed up computations, we used three folds for cross-fitting and used four-day mortality as our outcome of interest. Some observations had large weights $\mathsf{K}_{t,s}$ and $\mathsf{H}_{t,s}$, so we truncated the corresponding estimates at their 99% quantiles to avoid empirical violations of the positivity assumption. The results of the main analysis are presented in Table 1, where we contrast the mortality under an intervention to prevent censoring versus the mortality under an intervention to prevent censoring and to delay intubation by one day as specified above. Specifically, we estimate the total effect defined as $\mathsf{E}[Y(\bar{A}^{\mathbbm{d}},\bar{J}(\bar{A}^{\mathbbm{d}}))-Y(\bar{A},\bar{J}(\bar{A}))]$, where $\mathbbm{d}$ is defined in (8) and $Y$ is four-day survival rate under an intervention to prevent censoring. The indirect effect is thus defined as $\mathsf{E}[Y(\bar{A},\bar{J}(\bar{A}^{\mathbbm{d}}))-Y(\bar{A},\bar{J}(\bar{A}))]$, and the direct effect is $\mathsf{E}[Y(\bar{A}^{\mathbbm{d}},\bar{J}(\bar{A}^{\mathbbm{d}}))-Y(\bar{A},\bar{J}(\bar{A}^{\mathbbm{d}}))]$.
Effect Standard Error
Total 0.059 0.038
Direct (i.e., not through AKI) -0.024 0.039
Indirect (i.e., through AKI) 0.083 0.027
@@ 4.3 Results
Next, we aim to identify treatment effect modifiers that can be used to predict which patients are at a higher risk of a harmful effect through AKI, and which patients are likely to benefit due to all other mechanisms. Rigorously estimating interpretable conditional average treatment effects, or treatment effects given some covariate values, is notoriously difficult, especially when continuous covariates are considered (see for example Kennedy et al. (2017) for a review of the state-of-the-art in estimation of conditional effects). For simplicity and interpretability, we can approach this by following a strategy presented in Rudolph and Díaz (2022). Specifically, we regress the influence function (see Theorem 2) for the interventional direct effect of intubation onto baseline covariates. (It is easily shown that the conditional (on covariates) expectation of the (uncentered) influence function for the ATE in standard settings is equal to the conditional average treatment effect; for the extension of this principle to the LMTP case see equation 6 of Díaz et al. (2021). To derive the influence function of the direct and indirect effects, note that the influence function for the difference of two parameters is equal to the difference of the influence functions of those parameters.)
@@
Thus, we estimated univariate effect modification for each baseline variable by linear regression of the efficient influence function of the direct and indirect effect shown in equations 1 and 2 on each variable at a time, where we use the slope of the regression as a measure of effect modification. We show only the ten variables with the largest absolute value for univariate effect modification in Table 2. (If we were predominantly interested in the application of a treatment rule based on mediation effects, as described in Section 4.4, then it would be more appropriate to regress the influence function onto multiple covariates simultaneously; here we focus on the simple regressions to provide a more straightforward summary.)

### v3

@@ 6 Acknowledgments
The theory of semiparametric inference allows for the analysis of estimators of parameters (e.g., causal effects) without restricting $\mathsf{P}$ to a parametric model. Standard references include van der Vaart (1998) and Bickel et al. (1997). In our case, we are particularly interested in nonparametric inference which places no restrictions on $\mathsf{P}$ other than the identification assumptions, particularly those related to positivity.
@@
The requisite $n^{1/4}$-rate estimator convergence is achievable for some certain classes of flexible predictors, including LASSO (Bickel et al., 2009), highly adaptive LASSO (Benkeser and Van Der Laan, 2016), random forests (Wager and Walther, 2015), and neural networks (Chen and White, 1999) though we have not verified this fact for each estimator implemented in Super Learner. One may also consult the discussion of convergence rates for machine learning estimators in Section 3 of (Chernozhukov et al., 2018). We also note that if (correctly specified) parametric models are used, then the assumption can be easily verified. Overall, while the theoretical convergence rates of nonparametric estimators may be difficult to verify, we recommend the doubly robust procedure we have presented because it provides the best chance for appropriate uncertainty quantification while minimizing the possibility of model misspecification.
@@
$\mathsf{D}_{L,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{{}^{\prime}}_{t+1,s}\mathsf{H}_{t,s}\{\mathsf{Q}_{Z,s+1}-\mathsf{Q}_{L,s}\}+\sum_{s=t+1}^{\tau}\mathsf{K}^{{}^{\prime}}_{t+1,s}\mathsf{H}_{t,s-1}\{\mathsf{Q}_{L,s}-\mathsf{Q}_{Z,s}\}+\mathsf{Q}_{L,t}$ (S1)
$\mathsf{D}_{Z,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{{}^{\prime}}_{t,s}\mathsf{H}_{t,s}\{\mathsf{Q}_{Z,s+1}-\mathsf{Q}_{L,s}\}+\sum_{s=t}^{\tau}\mathsf{K}^{{}^{\prime}}_{t,s}\mathsf{H}_{t,s-1}\{\mathsf{Q}_{L,s}-\mathsf{Q}_{Z,s}\}+\mathsf{Q}_{Z,t}$ (S2)
$\mathsf{D}_{M,t}(\underaccent{\bar}{X}_{t},\underaccent{\bar}{m}_{t})=\sum_{s=t}^{\tau}\mathsf{K}^{*}_{t,s}[\prod_{k=t}^{s-1}\mathds{1}(M_{k}=m_{k})][\mathds{1}(M_{s}=m_{s})\mathsf{Q}_{M,s+1}-Q_{M,s}]+Q_{M,t}$ (S3)
@@
The efficient influence function for $\theta$ in the nonparametric model is given by
$\mathsf{S}(X,\eta)=\sum_{\bar{m}\in\bar{\mathcal{M}}}\left[\{\mathsf{D}_{Z,1}(X,\bar{m};\eta)-\varphi(\bar{m})\}\lambda(\bar{m})+\{\mathsf{D}_{M,1}(X,\bar{m};\eta)-\lambda(\bar{m})\}\varphi(\bar{m})\right]$
where $\bar{\mathcal{M}}$ is the space of possible values for the mediator variables.
Proof See the proof of Theorem 2 of Díaz et al. (2023b). ∎
@@
The algorithm used to estimate $\theta$ is given as Algorithm 1 in Díaz et al. (2023b), where in our case the IPW weights are replaced with probability ratios as described in Section 3.2. The algorithm consists of the following steps. First, we set $\hat{\mathsf{D}}_{Z,\tau+1}=Y$ and $\hat{\mathsf{D}}_{M,\tau+1}=1$. Then, beginning at $t=\tau$, we calculate $\hat{\mathsf{Q}}_{L,t}$ as the nonparametric regression prediction of $\hat{\mathsf{D}}_{Z,t+1}$, onto preceding variables. (The method is agnostic as to the choice of regression technique, but we recommend nonparametric or “data-adaptive” approaches; see Section 4.3 of the main text for the specific tools we use in our data analysis.) Then, we nonparametrically estimate the functions $\mathsf{g}$ and $\mathsf{G}$ (as in Section 3.2), and $\hat{\mathsf{D}}_{L,t}$ is calculated by plugging in all the nuisance functions estimated so far into its definition (in equation S1). Similarly, $\hat{\mathsf{Q}}_{Z,t}$ comes from a nonparametric regression of $\hat{\mathsf{D}}_{L,t}$, and $\hat{\mathsf{Q}}_{M,t}$ comes from a nonparametric regression of $\mathbbm{I}(M_{t}=m_{t})\hat{\mathsf{D}}_{M,t+1}$. Then $\hat{\mathsf{D}}_{Z,t}$ is estimated from equation S2 and $\hat{\mathsf{D}}_{M,t}$ is estimated from equation S3. This is repeated for $t=\tau-1,\tau-2,\dots,1$. Finally, the mean of $\hat{\mathsf{D}}_{Z,1}$ is used as an estimator for $\phi$, and the mean of $\hat{\mathsf{D}}_{M,1}$ is taken as an estimator of $\lambda$, which are plugged into the identifying formula of Theorem 1 to estimate $\theta$. (Note that the use of the $\mathsf{D}$-functions in this way is motivated by Lemma 1 of Díaz et al. (2023b).) The estimated variance of the efficient influence function (Theorem 1) is used to estimate the standard error. The procedure makes use of cross-fitting in all regression functions to ensure that the asymptotic linearity result above holds without restrictions on the complexity of the form of the machine learning estimators, thereby yielding valid statistical inference under mild conditions (Zheng and van der Laan, 2011; Chernozhukov et al., 2018). See Section 5 of Díaz et al. (2023b) for additional discussion of the convergence of the estimator described in that paper, which also applies to the current case.
S6 More on the structure of the observed data and intervention
S1 Measured covariates
@@
Baseline confounders include age, sex, race, ethnicity, body mass index (BMI), hospital location, home oxygen status, and comorbidities (e.g., hypertension, history of stroke, diabetes mellitus, coronary artery disease, active cancer, cirrhosis, asthma, chronic obstructive pulmonary disease, interstitial lung disease (ILD), HIV infection, and immunosuppression), and are included in $L_{1}$. Time-dependent confounders $Z_{t}$ and $L_{t}$ include vital signs (e.g., highest and lowest respiratory rate, oxygen saturation, temperature, heart rate, and blood pressure), laboratory results (e.g., C-Reactive Protein, BUN creatinine ratio, activated partial thromboplastin time, creatinine, lymphocytes, neutrophils, bilirubin, platelets, D-dimer, glucose, arterial partial pressure of oxygen, and arterial partial pressure of carbon dioxide).
S2 Missingness and censoring
In cases of missing baseline confounders, mean substitution with an additional variable to indicate missingness is used. For missing data at later timepoints, the last observation is carried forward. Patients are censored at their day of hospital discharge, as AKI and vital status were unknown after this point.
S3 Time of observation
Throughout, we have assumed that $\tau$, the length of follow-up, is equal for all observations. While this may sound like a restriction, we do not actually assume that all units are observed for the entire length of follow-up. Censored variables are all assigned a qualitative value indicating “censored,” and variables occurring after mortality are assigned a qualitative value indicating “mortality.” For our scientific study, we have $\tau=4$. For censored values, we adopt the procedure of including an intervention in our estimands to “prevent censoring,” which has the consequence that all outcome models at time $t$ are fitted only with units that are uncensored at time $t$.
S4 Censoring and survival
Although we are mainly scientifically interested in the intervention to delay intubation, we re-iterate here that part of the intervention we consider is to “prevent censoring”; this intervention applies to both treatment arms we are comparing (status quo and delay intubation). What this means practically is that outcomes are estimated for censored observations based on our estimates of the nuisance functions. While this may introduce some uncertainty relative to an uncensored data set, it does not induce bias unless the censoring at random assumption is violated. While censored individuals might be thought a priori to be, for example, more healthy (if they are discharged), our goal has been to control for enough baseline and time-varying covariates to make such conditional differences minimal. Statistically, our procedure has the effect that censored observations do not provide direct information about the effect of the intervention to delay intubation (because they do not provide outcome data), but they do inform the estimates of nuisance functions that contribute to the estimation of the effect.
However, we recognize that discharge could be considered a competing event and that incorporating that aspect of the problem could be a meaningful extension. Our analysis illustrates the method, but future work could extend this approach to incorporate competing risks, enhancing the interpretability of results in a substantive application. For a full-length discussion of the approach we are using for censoring, see Young et al. (2020).
S5 Survival analysis and effect parametrization
Often, “survival analysis” is performed by analyzing differences in hazard functions, either semi-parametrically through a proportional hazards assumption and Cox regression or through distribution-free hypothesis testing with the logrank test. Although such methods are plausible, in the context of causal inference it is more natural to define a specific parameter that quantifies the scientific effect to be estimated. In this case, the parameter is the difference in mortality rates between two different treatment regimes: one the status quo intervention, and one the intervention with delayed intubation. Thus, the form of the parameter is fundamentally no different from treatment effects in non-survival contexts except that censoring is considered. To add extra temporal dimensions to the results, one could repeat the analysis for multiple values of follow-up time or delay time, but we have chosen to estimate just one parameter for the current study.
S7 Effect modification in the AKI study
@@
Here, we aim to identify treatment effect modifiers that can be used to predict which patients are at a higher risk of a harmful effect through AKI, and which patients are likely to benefit due to all other mechanisms. Rigorously estimating interpretable conditional average treatment effects, or treatment effects given some covariate values, is notoriously difficult, especially when continuous covariates are considered (see for example Kennedy et al. (2017) for a review of the state-of-the-art in estimation of conditional effects). For simplicity and interpretability, we can approach this by following a strategy presented in Rudolph and Díaz (2022). Specifically, we regress the influence function (see Theorem 1) for the interventional direct effect of intubation onto baseline covariates. (It is easily shown that the conditional (on covariates) expectation of the (uncentered) influence function for the ATE in standard settings is equal to the conditional average treatment effect; for the extension of this principle to the LMTP case see equation 6 of Díaz et al. (2021). To derive the influence function of the direct and indirect effects, note that the influence function for the difference of two parameters is equal to the difference of the influence functions of those parameters.)
@@
Thus, we estimated univariate effect modification for each baseline variable by linear regression of the efficient influence function of the direct and indirect effect shown in equations 1 and 2 on each variable at a time, where we use the slope of the regression as a measure of effect modification. We show only the ten variables with the largest absolute value for univariate effect modification in Table S2. (If we were predominantly interested in the application of a treatment rule based on mediation effects, as described in Section 4.4, then it would be more appropriate to regress the influence function onto multiple covariates simultaneously; here we focus on the simple regressions to provide a more straightforward summary.)
First, we note that this analysis is exploratory; some subsets may have relatively small sample sizes and we have not calculated standard errors. However, interstitial lung disease (ILD) is the most important variable for effect modification for both the indirect effect through AKI and the direct effect through other mechanisms. The negative effect modification parameter for ILD in Table S2 implies that delaying IMV in patients with ILD has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients without ILD, whereas the positive effect modification parameter in Table S2 means that delaying IMV in patients with ILD has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without ILD. The former effect can be explained by the fact that delaying intubation may be riskier in patients with ILD; the latter effect (smaller in size) might be due to the fact that patients with ILD are more likely to experience mortality due to lung failure rather than AKI.
Likewise, Table S2 shows that delaying IMV in patients with cancer, chronic obstructive pulmonary disease (COPD), high lymphocyte count, costovertebral angle tenderness (CVA), coronary artery disease (CAD), high creatinine values, and those who used home oxygen, has a better effect on survival (i.e., is less harmful/more beneficial) due to AKI than in patients without those conditions. Delaying IMV in patients who have cirrhosis or who are current smokers has a worse effect on survival (i.e., is more harmful/less beneficial) due to AKI than in patients without those conditions.
The results in Table S2 show that delaying intubation in patients who are immunosuppressed, use home oxygen, or are white, has a worse effect on survival (i.e., is more harmful/less beneficial) due to mechanisms other than AKI than in patients outside those categories. Lastly, delaying intubation in patients with high creatinine values, asthma, CVA, cirrhosis, high bilirubin, and who are former smokers, has a better effect on survival (i.e., is less harmful/more beneficial) due to mechanisms other than AKI than in patients without those conditions.
Table S2: Most important effect modifiers for the direct and indirect effect of invasive mechanical ventilation on survival through acute kidney failure, as measured by the slope of a simple linear regression of the efficient influence function. Variables with negative slopes indicate populations with lower effects (calculated as intervention minus baseline; positive values indicate that delaying intubation is relatively less harmful/more beneficial).